OpenStudy (anonymous):

Find equations of the lines passing through (1, 3) and having the following characteristics. (a) Slope of -2/3 (b) Perpendicular to the line x + y = 0 (c) Passing through the point (2, 4) (d) Parallel to the x-axis

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

see when slope and one point is given: Use: \[y-y_1 = m(x - x_1)\] Can you plug in the values?? m = -2/3 and (x1, y1) = (1, 3) Can you do it??

OpenStudy (anonymous):

Hmm... but i don't understand the ab,c,d .. do I just get one final answer, or four?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

You have four different answers.. I am solving for a part..

OpenStudy (anonymous):

are you solving for part a?

OpenStudy (anonymous):

Yes i am..

OpenStudy (anonymous):

See I have provided you the formula can you plug in the value and solve??

OpenStudy (anonymous):

sure, let me try

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

y-3=-2/3(x-1)

OpenStudy (anonymous):

Yes.. Do you want to solve it further or not?? Or are their choices provided for this question??

OpenStudy (anonymous):

there aren't choices

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

So i don't think it matters. can u help with b,c, and d? :)

OpenStudy (anonymous):

If its perp, the slope should be 3/2

OpenStudy (anonymous):

See, equation of perpendicular line is given.. No, our first question is over.. You are using the slope of first part...

OpenStudy (anonymous):

We have a.. but we need b, c, and d.

OpenStudy (anonymous):

Oh right, never mind

OpenStudy (anonymous):

Wait dear let me explain I will make you clear each and everything..

OpenStudy (anonymous):

x+y=0

OpenStudy (anonymous):

Yes now can you find slope from this equation x = y = 0??

OpenStudy (anonymous):

*+

OpenStudy (anonymous):

Is it -1

OpenStudy (anonymous):

Yes you are right.. Now can you tell me the slope of the line that is perpendicular to it??

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yes.. Now m= 1 and (x1, y1) = (1,3) [given] Can you find its equation?? Same procedure..

OpenStudy (anonymous):

y-y1=1(x-x1)

OpenStudy (anonymous):

y-3=1(x-1)

OpenStudy (anonymous):

Yes now you are right that is B part..

OpenStudy (anonymous):

ok cool. now c

OpenStudy (anonymous):

passing thru 3,4

OpenStudy (anonymous):

y-4= m(x-3) now what slope goes in m?

OpenStudy (anonymous):

When two points are given we use Two point formula: \[\large \frac{y - y_1}{y_2 - y_1} = \frac{x-x_1}{x_2 - x_1}\]

OpenStudy (anonymous):

What values would go in that..

OpenStudy (anonymous):

(x1, y1) = (1,3) and (x2, y2) = (2, 4)

OpenStudy (anonymous):

I meant 2, not 3 !!

OpenStudy (anonymous):

Ok let me plug

OpenStudy (anonymous):

y-3/4-3=x-1/2-1 ??

OpenStudy (anonymous):

Ok take your time.. After plugging in use cross multiplication to solve further..

OpenStudy (anonymous):

y-3/1 = x-1/1 is that right so far.

OpenStudy (anonymous):

Yes you are going right..

OpenStudy (anonymous):

y-3=x-1

OpenStudy (anonymous):

y - 3 = x - 1 That is answer for C part..

OpenStudy (anonymous):

yay! okay lastly d

OpenStudy (anonymous):

parallel to x-axis

OpenStudy (anonymous):

D) see always remember the line parallel to x axis will have equation as : \(y = \pm c\) and line parallel to y axis has equation \(x = \pm c\) Here the required equation is: \(y = \pm c\) Getting??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

And one thing you must remember that: Slope of Line parallel to X-axis will be 0 and Slope of Line perpendicular to X-axis or Parallel to Y-axis will be Infinity.. So, here m = 0

OpenStudy (anonymous):

ok so we have y=+/-c, with the slope as 0...

OpenStudy (anonymous):